# Visual Anagrams Reveal Hidden Differences in Holistic Shape Processing Across Vision Models

### v2

@@ Abstract
Humans are able to recognize objects based on both local texture cues and the configuration of object parts, yet contemporary vision models primarily harvest local texture cues, yielding brittle, non-compositional features. Work on shape-vs-texture bias has pitted shape and texture representations in opposition, measuring shape relative to texture, ignoring the possibility that models (and humans) can simultaneously rely on both types of cues, and obscuring the absolute quality of both types of representation. We therefore recast shape evaluation as a matter of absolute configural competence, operationalized by the Configural Shape Score (CSS), which (i) measures the ability to recognize both images in Object-Anagram pairs that preserve local texture while permuting global part arrangement to depict different object categories. Across 86 convolutional, transformer, and hybrid models, CSS (ii) uncovers a broad spectrum of configural sensitivity with fully self-supervised and language-aligned transformers – exemplified by DINOv2, SigLIP2 and EVA-CLIP – occupying the top end of the CSS spectrum. Mechanistic probes reveal that (iii) high-CSS networks depend on long-range interactions: radius-controlled attention masks abolish performance showing a distinctive U-shaped integration profile, and representational-similarity analyses expose a mid-depth transition from local to global coding. A BagNet control, whose receptive fields straddle patch seams, remains at chance (iv), ruling out any “border-hacking” strategies. Finally, (v) we show that configural shape score also predicts other shape-dependent evals (e.g.,foreground bias, spectral and noise robustness). Overall, we propose that the path toward truly robust, generalizable, and human-like vision systems may not lie in forcing an artificial choice between shape and texture, but rather in architectural and learning frameworks that seamlessly integrate both local-texture and global configural shape. 111 Project Page: https://www.fenildoshi.com/configural-shape/ footnotetext: Preprint. Under Review.
@@ 1 Introduction
Human object recognition is remarkably robust: we can effortlessly identify objects across dramatic variations in texture, scale, viewpoint, and context because we can focus on aspects of global configuration that are stable across such local photometric quirks [1, 2, 3, 4, 5, 6]. By contrast, state-of-the-art vision networks still harvest local, high-frequency shortcuts [7, 8, 9]. This strategy achieves high ImageNet accuracy [10] but leaves models brittle under texture shifts, adversarial noise, and compositional out-of-distribution stresses [11, 12, 13, 14]. The failure arises because models often seize on spurious yet linearly separable features when multiple predictive cues are available [15, 16, 17]. These differences between models and humans are often studied using the shape–versus–texture bias diagnostic, which pits shape vs. texture using cue-conflict stimuli  [8], but this metric is inherently relative: scores rise whenever shape coding strengthens or when texture coding weakens, rendering the absolute fidelity of global shape ambiguous [18, 19, 20]. Effective vision systems should exploit both cues when helpful [21, 22, 23], motivating an absolute assessment of shape and texture processing A.1.
Figure 1: Object-Anagram task: a probe of configural shape perception. (A) visual-anagram example—an identical set of 16 square diffusion patches is spatially permuted to form two distinct objects, here a wolf and an elephant (one shared patch is outlined in red). (B) additional image pairs from the object-anagram benchmark. each pair comprises globally different objects built from the same unordered patch multiset, forcing any successful classifier to rely solely on the global arrangement of parts.
@@
By converting a long-standing theoretical critique into a falsifiable measurement and linking the resulting scores to identifiable computational mechanisms, this work advances the science of holistic shape perception and offers actionable design principles for future vision systems.
@@ 4 Vision Models
Collectively, this curated but diverse cohort will allow us to identify which architectural, training, and data regimes are predictive of high Configural Shape Score.
@@ 7 Disentangling the influence of object category and anagram puzzle pieces on configural shape score
We formalized these qualitative observations by computing two metrics (Fig. 4C) – Puzzle component influence and Category influence – over the last and penultimate layer of all models. Puzzle component influence is the difference in cosine similarity between pairs with identical puzzle pieces (anagram pairs) and those with different puzzle pieces and the matched category differences (different-parts/different-category). Category Influence is the average similarity advantage for same-category pairs over different-category pairs, irrespective of local puzzle pieces. We then measure whether the configural shape scores can predict these metrics across all the models. Fig. 5C shows this relationship for the last layer. Higher Configural Shape Score corresponded to lower Puzzle Component Influence (negative correlations: $r=-0.70$ at the last layer, $r=-0.57$ at the penultimate layer), suggesting that higher-CSS models focus less on details of the local part appearance. Conversely, Configural Shape Score correlated positively with Category Influence ($r=0.80$ at the last layer, $r=0.83$ at the penultimate layer), indicating that high-CSS models encode representations that are shared between images within a category, while discriminating between categories. This trend is observed even when considering representations from DINOv2 backbones, which are fully self-supervised and have no pressure to form abstract category representations ($r=0.94$ between CSS and Category Influence and $r=-0.86$ between CSS and Puzzle Component Influence).
@@ 9 From Configural Shape Score to Broad Shape-Dependent Performance
To what degree does having a high CSS predict other representational benefits and qualities? As shown in Fig. 5, we found that Configural Shape Score was significantly correlated with scores from several evals, including: 2) Robustness to Noise (r=0.81), testing each model’s performance across varying severity levels for five distinct noise types, as described in [11]. 2) Foreground-vs-Background Bias (r=0.76), tested using ImageNet-9 dataset from [68], quantifying the extent to which a model relies primarily on the foreground object rather than background information for classification. 3) Phase Dependence (r=0.73), assessed by swapping phase information in Fourier space between images and then measuring top-1 accuracy, quantifying the model’s reliance on phase information [69]. 4) Critical band masking strategy outlined by [14] (r=0.83), used to determine the bandwidth of spatial frequencies essential for accurate object recognition. In contrast, the shape-vs-texture bias score across these models showed weaker relationships to these evals (r=0.62 with Robustness to Noise; r=0.32 with Foreground-vs-Background Bias; r=0.52 with Phase Dependence and r=0.55 with Critical Band Masking). Statistical comparison using Williams’s test confirmed that CSS was a significantly better predictor of these metrics than shape-vs-texture bias (all p<0.001; see A.6 for test statistics and details). These results suggest that models with better configural shape scores also have other favorable and human-like perceptual qualities. See A.5 for more information about these evaluations, and A.7 and A.8 for feature attributions to qualitatively compare low- and high-CSS models on challenging stimuli from each benchmark.
@@ 10 Limitations and Discussion
Although the Object Anagram Dataset and the accompanying Configural Shape Score (CSS) provide a quantitative measure of holistic processing, several caveats warrant mention. First, the stimuli we generate are constrained by the priors of the diffusion model, and may explore only a subspace of configural encoding relationships. Second, this work targets whole-object configurations and therefore does not directly probe part-based compositionality, an orthogonal facet of shape reasoning that future work should address. Third, despite containing thousands of composites, the dataset is modest in scale compared with modern billion-image corpora, suggesting that larger or more ecologically varied stimuli could reveal subtler effects.
@@ A.1 Shape-vs-Texture Bias: A Useful but Incomplete Metric for Assessing Shape Representations
Shape-vs-Texture Bias(accuracy-corrected)=#Correct Shape Decisions#Correct (Shape + Texture)×#Correct Shape DecisionsTotal $\text{Shape-vs-Texture Bias}\\ \text{(accuracy-corrected)}\end{aligned}=\sqrt{\frac{\#\text{Correct Shape Decisions}}{\#\text{Correct (Shape + Texture)}}}\times\sqrt{\frac{\#\text{Correct Shape Decisions}}{\text{Total Trials}}}$
@@ A.4 Human Configural Shape score Estimate
A.5 Other Shape-dependent Evals
@@
Figure 7: Schematic of other Shape-dependent Evals
A.6 Statistical Comparison of Predictive Strength: CSS vs. Shape-vs-Texture Bias
@@
Eval Benchmark r(CSS, Eval Benchmark) r(Shape-vs-Texture Bias, Eval Benchmark) t-value p-value
@@
Foreground-vs-Background Bias 0.76 0.32 7.618 $<$0.0001
@@
Critical Band Masking 0.83 0.55 5.47 $<$0.0001
Table 3: Comparison of Configural Shape Score (CSS) and Shape-vs-Texture Bias
A.7 Feature Attribution for Challenging stimuli
@@
Figure 8: Feature Attribution Maps generated using Integrated Gradients for challenging stimuli - cue-conflict stimuli, phase swapped stimuli, visual anagrams
A.8 Feature Attributions of Anagrams in High-CSS model (DINOv2-B/14) are not Anagrams themselves

### v3

@@ Abstract
Humans are able to recognize objects based on both local texture cues and the configuration of object parts, yet contemporary vision models primarily harvest local texture cues, yielding brittle, non-compositional features. Work on shape-vs-texture bias has pitted shape and texture representations in opposition, measuring shape relative to texture, ignoring the possibility that models (and humans) can simultaneously rely on both types of cues, and obscuring the absolute quality of both types of representation. We therefore recast shape evaluation as a matter of absolute configural competence, operationalized by the Configural Shape Score (CSS), which (i) measures the ability to recognize both images in Object-Anagram pairs that preserve local texture while permuting global part arrangement to depict different object categories. Across 86 convolutional, transformer, and hybrid models, CSS (ii) uncovers a broad spectrum of configural sensitivity with fully self-supervised and language-aligned transformers – exemplified by DINOv2, SigLIP2 and EVA-CLIP – occupying the top end of the CSS spectrum. Mechanistic probes reveal that (iii) high-CSS networks depend on long-range interactions: radius-controlled attention masks abolish performance showing a distinctive U-shaped integration profile, and representational-similarity analyses expose a mid-depth transition from local to global coding. A BagNet control, whose receptive fields straddle patch seams, remains at chance (iv), ruling out any “border-hacking” strategies. Finally, (v) we show that configural shape score also predicts other shape-dependent evals (e.g.,foreground bias, spectral and noise robustness). Overall, we propose that the path toward truly robust, generalizable, and human-like vision systems may not lie in forcing an artificial choice between shape and texture, but rather in architectural and learning frameworks that seamlessly integrate both local-texture and global configural shape.  [^1] Project Page: https://www.fenildoshi.com/configural-shape/
@@ 1 Introduction
Human object recognition is remarkably robust: we can effortlessly identify objects across dramatic variations in texture, scale, viewpoint, and context because we can focus on aspects of global configuration that are stable across such local photometric quirks [1, 2, 3, 4, 5, 6]. By contrast, state-of-the-art vision networks still harvest local, high‑frequency shortcuts [7, 8, 9]. This strategy achieves high ImageNet accuracy [10] but leaves models brittle under texture shifts, adversarial noise, and compositional out‑of‑distribution stresses [11, 12, 13, 14]. The failure arises because models often seize on spurious yet linearly separable features when multiple predictive cues are available [15, 16, 17]. These differences between models and humans are often studied using the shape–versus–texture bias diagnostic, which pits shape vs. texture using cue‑conflict stimuli  [8], but this metric is inherently relative: scores rise whenever shape coding strengthens or when texture coding weakens, rendering the absolute fidelity of global shape ambiguous [18, 19, 20]. Effective vision systems should exploit both cues when helpful [21, 22, 23], motivating an absolute assessment of shape and texture processing A.1.
Figure 1: Object-Anagram task: a probe of configural shape perception. (A) visual‐anagram example—an identical set of 16 square diffusion patches is spatially permuted to form two distinct objects, here a wolf and an elephant (one shared patch is outlined in red). (B) additional image pairs from the object‐anagram benchmark. each pair comprises globally different objects built from the same unordered patch multiset, forcing any successful classifier to rely solely on the global arrangement of parts.
@@
By converting a long‑standing theoretical critique into a falsifiable measurement and linking the resulting scores to identifiable computational mechanisms, this work advances the science of holistic shape perception and offers actionable design principles for future vision systems.
@@ 4 Vision Models
Collectively, this curated but diverse cohort will allow us to identify which architectural, training, and data regimes are predictive of high Configural Shape Score (see A.11 for the complete list of all the evaluated models)..
@@ 6 Long-range Contextual Interactions lead to higher Configural Shape Scores in Vision Transformers
To further test the critical role of these intermediate layers, especially in deeper architectures where single-block ablations can be too subtle, we designed a more stringent cumulative ablation experiment on the full DINOv2 family. In this analysis, we measured the irrecoverable contribution of intermediate layers by systematically ablating long-range interactions in all blocks before or after a selected layer, and measuring it’s impact on the class token’s representation from the last block. The full methodology and results are detailed in A.5. The biggest drop was seen in the intermediate blocks across all model sizes, confirming that the mid-depth layers were indeed the locus of the most critical computations for configural processing in these models.
7 Relational Positional Encodings Boost Configural Shape Scores
Table 1: Configural Shape Scores (CSS) for ViT-B/16 variants.
Model Variant (ViT-B/16) CSS (%)
Relative Positional Embeddings 38.89
RoPE (Rotary Position Embedding) 51.38
RoPE (Rotary Position Embedding) + Absolute PE 51.38
Mixed RoPE 48.61
Mixed RoPE + Absolute PE 52.77
While long-range attention is necessary for configural processing, the underlying mechanism for encoding spatial positions is also critical. To isolate this factor, we evaluated five variants of the supervised ViT-B/16 architecture, each with a different positional encoding (PE) scheme (Table 1). While the standard model using learned absolute PEs achieved a CSS of 23.61%, models equipped with relative PEs, such as Rotary Position Embeddings (RoPE), more than doubled this score to over 51%. This result demonstrates that a model’s ability to process the relational spatial structure of an image is criticial for holistic shape perception, and that CSS rewards this capacity.
@@ 8 Disentangling the influence of object category and anagram puzzle pieces on configural shape score
We formalized these qualitative observations by computing two metrics (Fig. 4C) – Puzzle component influence and Category influence – over the last and penultimate layer of all models. Puzzle component influence is the difference in cosine similarity between pairs with identical puzzle pieces (anagram pairs) and those with different puzzle pieces and the matched category differences (different-parts/different-category). Category Influence is the average similarity advantage for same-category pairs over different-category pairs, irrespective of local puzzle pieces. We then measure whether the configural shape scores can predict these metrics across all the models. Fig. 5C shows this relationship for the last layer. Higher Configural Shape Score corresponded to lower Puzzle Component Influence (negative correlations: r=0.70r=-0.70 at the last layer, r=0.57r=-0.57 at the penultimate layer), suggesting that higher-CSS models focus less on details of the local part appearance. Conversely, Configural Shape Score correlated positively with Category Influence ($r=0.80$ at the last layer, $r=0.83$ at the penultimate layer), indicating that high-CSS models encode representations that are shared between images within a category, while discriminating between categories. This trend is observed even when considering representations from DINOv2 backbones, which are fully self-supervised and have no pressure to form abstract category representations ($r=0.94$ between CSS and Category Influence and r=0.86r=-0.86 between CSS and Puzzle Component Influence).
@@ 10 From Configural Shape Score to Broad Shape-Dependent Performance
To what degree does having a high CSS predict other representational benefits and qualities? As shown in Fig. 5, we found that Configural Shape Score was significantly correlated with scores from several evals, including: 1) Robustness to Noise (r=0.81), testing each model’s performance across varying severity levels for five distinct noise types, as described in [11]. 2) Foreground-vs-Background Bias (r=0.76), tested using ImageNet-9 dataset from [68], quantifying the extent to which a model relies primarily on the foreground object rather than background information for classification. 3) Phase Dependence (r=0.73), assessed by swapping phase information in Fourier space between images and then measuring top-1 accuracy, quantifying the model’s reliance on phase information [69]. 4) Critical band masking strategy outlined by [14] (r=0.83), used to determine the bandwidth of spatial frequencies essential for accurate object recognition. In contrast, the shape-vs-texture bias score across these models showed weaker relationships to these evals (r=0.62 with Robustness to Noise; r=0.32 with Foreground-vs-Background Bias; r=0.52 with Phase Dependence and r=0.55 with Critical Band Masking). Statistical comparison using Williams’s test confirmed that CSS was a significantly better predictor of these metrics than shape-vs-texture bias (all p<0.001; see A.7 for test statistics and details). These results suggest that models with better configural shape scores also have other favorable and human-like perceptual qualities. See A.6 for more information about these evaluations, and A.8 and A.9 for feature attributions to qualitatively compare low- and high-CSS models on challenging stimuli from each benchmark.
@@ 11 Limitations and Discussion
Although the Object Anagram Dataset and the accompanying Configural Shape Score (CSS) provide a quantitative measure of holistic processing, several caveats warrant mention. First, the stimuli we generate are constrained by the priors of the diffusion model, and may explore only a subspace of configural encoding relationships. Furthermore, we use a uniform "black paint texture" to ensure local cues are perfectly matched, though this raises the possibility of out-of-distribution effects. However, our analyses show that even models with poor CSS can classify single anagram images successfully, suggesting the primary failure is in configural processing, not in handling the texture itself. Second, this work targets whole-object configurations and therefore does not directly probe part-based compositionality, an orthogonal facet of shape reasoning that future work should address. Third, despite containing thousands of composites, the dataset is modest in scale compared with modern billion-image corpora, suggesting that larger or more ecologically varied stimuli could reveal subtler effects. To that end, we created a 20x larger set of 1440 anagram pairs to test the robustness of our main findings. Re-evaluating all 86 models confirmed that the CSS metric is highly stable: the relative model rankings were preserved, and the scores from the original and expanded datasets showed a correlation of r=0.99 (see Appendix A.10).
@@
Our results suggest that high configural competence stems primarily from training objectives that enforce local-to-global consistency. For instance, DINOv2’s teacher-student objective, which trains a student network on local image crops to match the output distribution of a teacher network seeing global crops, directly rewards the integration of long-range information. This view-consistency objective leads to a more powerful signal than standard data augmentations alone or purely reconstructive approaches like MAE, whose local pixel-prediction task is insufficient on its own to produce the mid-layer “configural flip” characteristic of high-CSS models. Language supervision (BEiT-v2, EVA-CLIP, CLIP, SigLIP) would also enforce alignment between standard local-image crops and a global text caption, and this form of local-global consistency objective could also place an emphasis on multiple distant concepts in the image. Together, these findings point to a clear roadmap for building more shape-aware systems, which should prioritize (1) global-local consistency objectives, paired with (2) architectures capable of long-range integration, and (3) evaluated with explicit shape metrics like CSS.
@@ A.1 Shape-vs-Texture Bias: A Useful but Incomplete Metric for Assessing Shape Representations
[…]
@@ A.4 Human Configural Shape score Estimate
A.5 Cumulative Ablation Analysis on DINOv2 Models
To more precisely characterize the role of intermediate layers for configural processing, we performed a cumulative ablation study on the DINOv2 model family (S/14, B/14, L/14, and G/14) and the ViT-B/16 baseline. We hypothesized that the single-block ablation presented in Section 6 might be too subtle to reveal the full processing dynamics in deeper models. To address this, we designed an ablation experiment with two conditions: (1) ablating all blocks up to block ’n’ to isolate the contribution of the network up to that point, and (2) ablating all blocks after block ’n’ to test for recovery by later layers. The ablation follows the same strategy of masking the attention in the targeted blocks using a mask window of 2 nearby patches for all the patches, hence blocking long-range interactions.
Fig. 7 shows the effect of these ablations on the final class token representation, measured by its cosine similarity to the token from the original, unablated model. In the “Ablate all before” condition (blue line), the similarity starts high and drops steeply as intermediate blocks are removed, while the “Ablate all after” condition (orange line) shows the opposite trend, creating the characteristic crossover point in the intermediate layers of each model. The results demonstrate that across all DINO models tested, the drop in CSS is highest and most irrecoverable in the intermediate layers. This confirms that even for the large and giant DINO variants, the mid-depth layers are the critical locus for integrating local features into a global, configural shape representation.
Figure 7: Effect of cumulative ablation on class token representation. The y-axis shows the cosine similarity between the final class token of the original model and the ablated model. The blue line (Ablate all before) shows the condition when ablating long-range interactions that is ‘accumulated’ in all blocks before block n. The orange line (Ablate all after) shows the condition when ablating long-range interactions that could be ‘recovered’ in all blocks after block n.
A.6 Other Shape-dependent Evals
@@
Figure 8: Schematic of other Shape-dependent Evals.
A.7 Statistical Comparison of Predictive Strength: CSS vs. Shape-vs-Texture Bias
@@
Eval Benchmark r(CSS, Eval) r(Shape-vs-Texture Bias, Eval) t-value p-value
@@
Foreground-vs-Background Bias 0.76 0.32 7.618 <<0.0001
@@
Critical Band Masking 0.83 0.55 5.47 <<0.0001
Table 4: Comparison of Configural Shape Score (CSS) and Shape-vs-Texture Bias.
A.8 Feature Attribution for Challenging stimuli
@@
Figure 9: Feature Attribution Maps for challenging stimuli. Maps generated using Integrated Gradients for cue-conflict stimuli, phase swapped stimuli, and visual anagrams.
A.9 Feature Attributions of Anagrams in High-CSS model (DINOv2-B/14) are not Anagrams themselves